OpenStudy (anonymous):

Verify the function and state its domain. Indicate which theorems you are using, and which functions you are assuming to be continuous: 1/(sqrt{x+2})

OpenStudy (anonymous):

Will the domain be (0, infinity) ?

OpenStudy (anonymous):

x+2>0 x>-2 ????

OpenStudy (anonymous):

\[\sqrt{x+2}\] it is the denominator, which means that it can not be 0. and \[\sqrt{x+2}\]means X+2>=0 so x+2>0 so X >-2

OpenStudy (anonymous):

Okay, how do I prove it is continuous?